# On $[1,2]$-Domination in Interval and Circle Graphs

### v2

@@ Abstract
A subset $S$ of vertices in a graph $G=(V,E)$ is Dominating Set if each vertex in $V(G)\setminus S$ is adjacent to at least one vertex in $S$. Chellali et al. in 2013, by restricting the number of neighbors in $S$ of a vertex outside $S$, introduced the concept of $[1,j]$-dominating set. A set $D\subseteq V$ of a graph $G=(V,E)$ is called $[1,j]$-Dominating Set of $G$ if every vertex not in $D$ has at least one neighbor and at most $j$ neighbors in $D$. The Minimum $[1,j]$-Domination problem is the problem of finding the minimum set $D$. Given a positive integer $k$ and a graph $G=(V,E)$, the $[1,j]$-Domination Decision problem is to decide whether $G$ has $[1,j]$-dominating set of cardinality at most $k$. A polynomial-time algorithm was obtained in split graphs for a constant $j$ in contrast to the classic Dominating Set problem which is NP-hard in split graphs. This result motivates us to investigate the effect of restriction $j$ on the complexity of $[1,j]$-domination problem on various classes of graphs. Although for $j\geq 3$, it has been proved that the minimum of classical domination is equal to minimum $[1,j]$-domination in interval graphs, the complexity of finding the minimum $[1,2]$-domination in interval graphs is still outstanding. In this paper, we propose a polynomial-time algorithm for computing a minimum $[1,2]$ on interval graphs by a dynamic programming technique. Next, on the negative side, we show that the minimum $[1,2]$-dominating set problem on circle graphs is $NP$-complete.
@@ 1 Introduction
One of the active research areas in graph theory is the study of the domination problem and its variant. Applying some restrictions on the domination problem leads to various extensions of the domination problem such as independent domination (Extra restriction is: each vertex in dominating set, like $D$, is not adjacent to a vertex in $D$), total domination (Extra restriction is: each vertex in $D$ is adjacent to at least one vertex in $D$) and, $[1,j]$-domination (Extra restriction is: each vertex not in $D$ is adjacent to at least one and at most $j$ vertices in $D$). Combinatorial and algorithmic results have been widely obtained on the various extensions of domination problems. A discussion of these results can be found in [9].
@@ 1.1 Definition and notation
Let $G=(V,E)$ is a simple graph with vertex set $V$ and edge set $E$. The order of graph $G$ is defined as $n=|V(G)|$. The open neighborhood of a vertex $v\in V$, denoted by $N(v)$, is defined as $\{u:\{u,v\}\in E\}$. Also the closed neighborhood of $v$ is defined as $N[v]=N(v)\cup\{v\}$. We use $N(S)$ and $N[S]$ to denote the open and closed neighborhood $S$, for a set $S\subseteq V$, respectively. That is, $N[S]=\bigcup_{v\in S}N[v]$ and $N(S)=N[S]\setminus S$. The complete bipartite graph $K_{1;3}$ is called claw and, a graph without a claw as an induced subgraph is called claw-free.
@@
A set $S\subseteq V$ is a dominating set if every vertex not in $S$ is adjacent to at least one vertex in $S$. The cardinality of the smallest dominating set, denoted by $\gamma(G)$, is called domination number. Finding a minimum dominating set was one of the first problems shown NP-hard in [6]. A set $D\subseteq V$ is called a $[1,j]$-dominating set of $G$ if for each $v\in V\setminus D$ we have $1\leq|N(v)\cap D|\leq j$, i.e. $v$ is adjacent to at least one but not more than $j$ vertices in $D$. Every graph has at least one $[1,j]$-dominating set since the set of vertices $V$ is itself a $[1,j]$-dominating set. The size of the smallest $[1,j]$-dominating sets of $G$ is denoted by $\gamma[1,j](G)$. For special case, where $j=2$, this domination is already known as Quasiperfect Domination. In the decision version of the $[1,j]$-dominating set problem, the input is a graph $G$ and a positive integer $k$ and, the objective is to test whether there is a $[1,j]$-dominating set of size at most $k$.
@@
Let $I=\{I_{1},I_{2},\dots,I_{n}\}$ be a set of intervals on real line. A graph $G=(V,E)$ is called an interval graph if each of its vertices can be associated with an interval in $I$ and two vertices are adjacent if and only if the corresponding intervals have intersection. The set $I$ is called the interval representation of an interval graph $G$. The intervals in $I$ are indexed by increasing the right endpoints. Figure 1 shows an interval graph and its internal representation.
@@
A linear time algorithm proposed in [2] to recognize an interval graph and the corresponding intervals of the vertices of an interval graph. The induced subgraphs of interval graphs are also interval [8] which is called hereditary property of the interval graphs. A proper interval graph is an interval graph such that no interval is completely contained within another interval. If all the intervals have the same length, its corresponding interval graph called a unit interval graph.
@@
A circle graph is a graph $G=(V,E)$ such that there is a one-to-one mapping between vertices in $V$ and a set $C$ of chords in a circle. Two vertices in $V$ are adjacent if and only if the corresponding chords in $C$ intersect. $C$ is called the chord intersection model for $G$. Equivalently, the vertices of a circle graph can be placed in the one-to-one correspondence the intervals such that two vertices are adjacent if and only if the corresponding intervals overlap, but neither contains the other. $I$ is called the interval representation of the circle graph. There exists a polynomial transformation between a representation of a circle graph, the set of its chords or its intervals representation. Moreover those are equivalent.(See Figure 2)
Figure 2: (a) A circle graph $G$ on $8$ vertices (b)The circle representation of $G$ (c)The interval representation of $G$
@@ 1.2 Short review of $[1,j]$-Domination
Chellali et al. in [3] introduced the concept of $[1,j]$-dominating set with name $(1,j)$-set, Although this concept for the constant $j=2$ already known as quasiperfect domination and defined by Dejter [4]. Also, $[1,j]$-dominating set fits into the general framework of $(\rho,\sigma)$-sets of graphs, which was first introduced by Telle in [17]. A $(\rho,\sigma)$-set is a set $D_{(\rho,\sigma)}\subseteq V$ such that for every vertex $v\in V$, we have $|N(v)\cap D_{(\rho,\sigma)}|\in\rho$ if $v\in D_{(\rho,\sigma)}$, and $|N(v)\cap D_{(\rho,\sigma)}|\in\sigma$ if $v\notin D_{(\rho,\sigma)}$. Discussion about the complexity of $[1,j]$-dominating set in various graph have been done extensively. The authors in [3] raised several open problems. One major open question is in what classes of graphs the $\gamma[1,j](G)$ is equal to $\gamma(G)$. In [3], it was shown that in $P_{4}$-free graphs, claw-free graphs, and caterpillars $\gamma[1,j](G)$ is equal to $\gamma(G)$. In [16] it was shown that $\gamma[1,j](G)=\gamma(G)$ when $G$ is a grid graph. It was proved in [19], that the value of $\gamma_{[1,2]}(G)$ in the classes of planar, bipartite, and triangle-free graphs can be equal to the entire set of vertices. So, $\gamma_{[1,2]}$ necessarily is not equal to $\gamma$ in these classes. In [5], the authors showed that in the interval graph and permutation graph $\gamma_{[1,3]}(G)=\gamma(G)$. Moreover, it was shown that there exists an interval graph $H$ for which $\gamma(H)<\gamma_{[1,2]}(H)$. Although they prove that for every unit interval graph $G$ we have $\gamma_{[1,2]}(G)=\gamma(G)$, the complexity of finding $\gamma_{[1,2]}$ in non-proper interval graphs remains unsolved.
@@
Another open question in [3] was in which graph classes $[1,j]$-dominating set is efficiently solvable. On the negative side, it was also proven that the problem is $NP$-complete even for bipartite graphs[3]. In [1], the $[1,j]$-dominating sets problem was shown to be $NP$-hard even for chordal and planar graphs. In [5], the authors showed the complexity of the decision problem of whether the $\gamma_{[1,j]}(G)=\gamma(G)$ is an $NP$ problem. On the positive side, The linear-time algorithms proposed for computing $\gamma_{[1,2]}(G)$ when $G$ is a tree [7], $G$ is a block graph [14] and $G$ is a serial-parallel graph [15]. For a constant $j$, a polynomial-time algorithm running in roughly $O(n^{j}p(\log n))$ where $p$ is a polynomial function, was obtained for $n$-vertex split graphs [1]. This is in contrast to the classic dominating set problem which is NP-hard for this class of graphs. Moreover in [10], a lower bound for computing $[1,j]$-dominating set in the split graph shown, i.e. there is no algorithm with running time $O(n^{j-\epsilon})$ where $\epsilon>0$ unless $P=NP$. The complexity of $[1,j]$-dominating set problem in various graph classes is summarized in Figure 3. It should be noted that this summary is far from being comprehensive and several other classes of graphs exist. In this paper, we focus only on the circle and interval graphs.
@@
Figure 3: Inclusion relations among well-studied classes of graphs [18] -Complexity of $[1,j]$-dominating set restricted to various classes of graphs in a hierarchy of graphs - $NPc$ indicate that the problem is $NP$-complete, $P$ indicates the problem has polynomial algorithm, and question mark indicate open problem.
@@ 2 Algorithm for interval graphs
Roberts [13] has shown that an interval graph is proper if and only if it contains no induced copy of $K_{1;3}$. In [4], it was shown that for any $K_{1;3}$-free graph $\gamma(G)=\gamma_{[1,2]}(G)$. So, easily the following lemma is concluded.
Lemma 1.
@@
Another proof for the above lemma was shown in [5]. The authors also show that for every interval graph $G$, $\gamma(G)=\gamma_{[1,3]}(G)$ and, there exists graph $G$ such that $\gamma(G)\leq\gamma_{[1,2]}(G)$, but the complexity of finding $[1,2]$-dominating set in interval graphs remained open.
In this section, we propose an $O(n^{4})$-time algorithm for computing $\gamma_{[1,2]}(G)$, where $G$ is an interval graph of order $n$. To present and prove the correctness of our algorithm, we need the following lemmas and notation. Let $G=(V,E)$ be an interval graph. Ramalingam and Rangan [12] proposed a numbering of vertices of $G$; they stated the following result.
Corollary 1 (Ramalingam and Rangan [12]).
For each interval graph $G=(V,E)$ of order $n$, there is a numbering $(1,2,\ldots,n)$ of vertices in $V$ such that if $ik\in E$ and $i<j<k$, then $jk\in E$.
This numbering of vertices in a given interval graph $G=(V,E)$ can be computed in $O(|V|+|E|)$ time [12]. Let $a$ and $b$ be integers such that $1\leq a\leq b\leq n$. We have to introduce the following notation.
@@
    $\gamma^{0}_{[1,2]}(i)=\min\{|D|:D$ is an $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i\notin D\}$,
@@
    $\gamma^{0}_{[1,2]}(j,i)=\min\{|D|:D$ is an $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $x\notin D$ for all $x\in[j,i]\}$,
@@
    $\gamma^{0}_{[1,2]}(j,i:k^{\prime})=\min\{|D|:D$ is an $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $k^{\prime}\in D$ and $x\notin D$ for all $x\in[j,i]\setminus\{k^{\prime}\}\}$, where $k^{\prime}\in[j,i)$,
@@
    $\gamma^{1}_{[1,2]}(i)=\min\{|D|:D$ is an $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i\in D\}$,
@@
    $\gamma^{1}_{[1,2]}(j,i:i)=\min\{|D|:D$ is an $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i\in D$ and $x\notin D$ for all $x\in[j,i)\}$,
@@
    $\gamma^{1}_{[1,2]}(k,i:i,j)=\min\{|D|:D$ is an $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i,j\in D$ and $x\notin D$ for all $x\in[k,i)\setminus\{j\}\}$,
@@
    $\gamma^{1}_{[1,2]}(l,i:i,j,k)=\min\{|D|:D$ is an $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i,j,k\in D$ and $x\notin D$ for all $x\in[l,i)\setminus\{j,k\}\}$,
@@
    $\gamma^{11}_{[1,2]}(l,i:i,j,k)=\min\{|D|:D$ is an $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i,j,x\in D$ for all $x\in[l,k]$ and $y\notin D$ for all $y\in(k,i)\setminus\{j\}\}$.
@@
Figure 4: (a) a set of intervals corresponding to the vertex set of the interval graph $G$ (b) a numbering $(1,2,\ldots,10)$ of vertices of $G$ satisfying the condition of Corollary 1, and (c) values $\mathtt{low}(i)$ and $\mathtt{maxlow}(i)$ for all $i\in[1,10]$
In the rest of this section, we assume that $G=(V,E)$ is an interval graph of order $n$, and without loss of generality with a numbering $(1,2,\ldots,n)$ of vertices of $G$ satisfying the condition of Corollary 1.
Corollary 2 (Poureidi [11]).
@@
Proof.
It is clear that the case (i) holds. Now, we prove the case (ii). Let $D$ be an $[1,2]$-dominating set of $G[1,i]$ with minimum cardinality such that $a\notin D$ for all $a\in[i^{\prime},i]$. So, $|D|=\gamma^{0}_{[1,2]}(i^{\prime},i)$. Since $i\notin D$, at least one and at most two of vertices adjacent to $i$ are in $D$. If $i^{\prime}\leq\mathtt{low}(i)$, then there is no vertex in $D$ to dominate $i$, that is, $D$ is not an $[1,2]$-dominating set of $G[1,i]$. Hence, $\mathtt{low}(i)<i^{\prime}$. By the case (ii) of Corollary 2, there is a vertex $x\in[\mathtt{maxlow}(i),i]$ such that $xy\notin E$ for each vertex $y\in[1,\mathtt{maxlow}(i))$. Thus, if $a\notin D$ for all $a\in[\mathtt{maxlow}(i),i]$, then $D$ is not an $[1,2]$-dominating set of $G[1,i]$ and so at least one of vertices of $[\mathtt{maxlow}(i),i)$ is in $D$. Since $a\notin D$ for all $a\in[i^{\prime},i]$, we obtain that $\mathtt{maxlow}(i)<i^{\prime}$ and at least one of vertices of $[\mathtt{maxlow}(i),i^{\prime})$ is in $D$. Assume that $j\in[\mathtt{maxlow}(i),i^{\prime})$ is a vertex in $D$ such that $a\notin D$ for all $a\in(j,i]$. We obtain that $D$ is an $[1,2]$-dominating set of $G[1,j]$. Let $b\in(j,i]$ be a vertex such that $\mathtt{low}(b)\leq\mathtt{low}(a)$ for all $a\in(j,i]$, that is, $\mathtt{low}(b)=\mathtt{low}(j+1,i)$. See Figure 5. Clearly, $\mathtt{low}(j+1,i)\leq\mathtt{low}(i)$. Because $b\notin D$ and $j\in D$, at most one of vertices of $[\mathtt{low}(b),j)$ is in $D$. Hence, either no vertex of $[\mathtt{low}(b),j)$ is in $D$, see Figure 5(a), or exactly one vertex of $[\mathtt{low}(b),j)$ is in $D$, see Figure 5(b). In the following we consider these two cases.

### v4

@@ Abstract
A subset $S$ of vertices in a graph $G=(V,E)$ is a Dominating Set if each vertex in $V(G)\setminus S$ is adjacent to at least one vertex in $S$. Chellali et al. in 2013, by restricting the number of neighbors in $S$ of a vertex outside $S$, introduced the concept of $[1,j]$-dominating set. A set $D\subseteq V$ of a graph $G=(V,E)$ is called a $[1,j]$-Dominating Set of $G$ if every vertex not in $D$ has at least one neighbor and at most $j$ neighbors in $D$. The Minimum $[1,j]$-Domination problem is the problem of finding the minimum $[1,j]$-dominating set $D$. Given a positive integer $k$ and a graph $G=(V,E)$, the $[1,j]$-Domination Decision problem is to decide whether $G$ has a $[1,j]$-dominating set of cardinality at most $k$. A polynomial-time algorithm was obtained in split graphs for a constant $j$ in contrast to the Dominating Set problem which is NP-hard for split graphs. This result motivates us to investigate the effect of restriction $j$ on the complexity of $[1,j]$-domination problem on various classes of graphs. Although for $j\geq 3$, it has been proved that the minimum of domination is equal to minimum $[1,j]$-domination in interval graphs, the complexity of finding the minimum $[1,2]$-domination in interval graphs is still outstanding. In this paper, we propose a polynomial-time algorithm for computing a minimum $[1,2]$-dominating set on interval graphs by a dynamic programming technique. Next, on the negative side, we show that the minimum $[1,2]$-dominating set problem on circle graphs is $NP$-complete.
@@ 1 Introduction
One of the active research areas in graph theory is the study of the domination problem and its variants. Applying specific restrictions on the domination problem leads to various extensions of the domination problem. In independent domination, the additional restriction is that every vertex in dominating set, like $D$, is not adjacent to any other vertex in $D$. In total domination, the additional restriction is that every vertex in $D$ is adjacent to at least one vertex in $D$. In $[1,j]$-domination, the additional restriction is that every vertex not in $D$ is adjacent to at least one and at most $j$ vertices in $D$. Combinatorial and algorithmic results have been widely obtained on the various extensions of domination problems. A discussion of these results can be found in Haynes et al. (1998).
@@ 1.1 Definition and notation
Let $G=(V,E)$ be a simple graph with vertex set $V$ and edge set $E$. The order of the graph $G$ is defined as $n=|V(G)|$. The open neighborhood of a vertex $v\in V$, denoted by $N(v)$, is defined as $\{u:\{u,v\}\in E\}$. Also, the closed neighborhood of $v$ is defined as $N[v]=N(v)\cup\{v\}$. We use $N(S)$ and $N[S]$ to denote the open and closed neighborhood $S$, for a set $S\subseteq V$, respectively. That is, $N[S]=\bigcup_{v\in S}N[v]$ and $N(S)=N[S]\setminus S$. The complete bipartite graph $K_{1;3}$ is called a claw, and a graph without a claw as an induced subgraph is called claw-free.
@@
A set $S\subseteq V$ is a dominating set if every vertex not in $S$ is adjacent to at least one vertex in $S$. The cardinality of the smallest dominating set, denoted by $\gamma(G)$, is called the domination number. Finding a minimum dominating set was one of the first problems shown to be NP-hard in Garey and Johnson (1979). A set $D\subseteq V$ is called a $[1,j]$-dominating set of $G$ if for each $v\in V\setminus D$ we have $1\leq|N(v)\cap D|\leq j$, i.e. $v$ is adjacent to at least one but not more than $j$ vertices in $D$. Every graph has at least one $[1,j]$-dominating set since the set of all vertices $V$ is itself a $[1,j]$-dominating set. The size of the smallest $[1,j]$-dominating set of $G$ is denoted by $\gamma[1,j](G)$. In the special case where $j=2$, this domination is already known as Quasiperfect Domination. In the decision version of the $[1,j]$-dominating set problem, the input is a graph $G$ and a positive integer $k$ and the objective is to test whether there is a $[1,j]$-dominating set of size at most $k$.
@@
Let $I=\{I_{1},I_{2},\dots,I_{n}\}$ be a set of intervals on the real line. A graph $G=(V,E)$ is called an interval graph if each of its vertices can be associated with an interval in $I$ and two vertices are adjacent if and only if the corresponding intervals intersect. The set $I$ is called the interval representation of the interval graph $G$. The intervals in $I$ are indexed by increasing the right endpoint. Figure 1 shows an interval graph and its interval representation.
@@
A linear-time algorithm was proposed by Booth and Lueker (1976) to recognize an interval graph and the corresponding intervals of the vertices of an interval graph. The induced subgraphs of interval graphs are also interval graphs which is called hereditary property of the interval graphs Golumbic (1980). A proper interval graph is an interval graph in which no interval is completely contained within another interval. If all the intervals have the same length, the corresponding interval graph is called a unit interval graph.
@@
A circle graph is a graph $G=(V,E)$ such that there is a one-to-one mapping between the vertices in $V$ and a set $C$ of chords in a circle. Two vertices in $V$ are adjacent if and only if the corresponding chords in $C$ intersect. $C$ is called the chord intersection model for $G$. Equivalently, the vertices of a circle graph can be placed in one-to-one correspondence with intervals such that two vertices are adjacent if and only if the corresponding intervals overlap, but neither contains the other. $I$ is called the interval representation of the circle graph. There exists a polynomial transformation between a representation of a circle graph, either the set of its chords or its interval representation. (See Figure 2)
Figure 2: (a) The circle graph $G$ on $8$ vertices (b)The circle representation of $G$ (c)The interval representation of $G$
@@ 1.2 Short review of $[1,j]$-Domination
Chellali et al. (2013) introduced the concept of a $[1,j]$-dominating set with name $(1,j)$-set. Although this concept for the constant $j=2$ was already known as quasiperfect domination and defined by Dejter (2009). Also, the $[1,j]$-dominating set fits into the general framework of $(\rho,\sigma)$-sets of graphs, which was first introduced in Telle (1994). A $(\rho,\sigma)$-set is a set $D_{(\rho,\sigma)}\subseteq V$ such that for every vertex $v\in V$, we have $|N(v)\cap D_{(\rho,\sigma)}|\in\rho$ if $v\in D_{(\rho,\sigma)}$, and $|N(v)\cap D_{(\rho,\sigma)}|\in\sigma$ if $v\notin D_{(\rho,\sigma)}$. Discussion about the complexity of $[1,j]$-dominating set in various graphs has been done extensively. Chellali et al. (2013) raised several open problems. One major open question is in what classes of graphs the $\gamma[1,j](G)$ is equal to $\gamma(G)$. Chellali et al. (2013) showed that in $P_{4}$-free graphs, claw-free graphs, and caterpillars $\gamma[1,j](G)$ is equal to $\gamma(G)$. Sharifani et al. (2020b) showed that $\gamma[1,j](G)=\gamma(G)$ when $G$ is a grid graph. It was proved in Yang and Wu (2014), that the value of $\gamma_{[1,2]}(G)$ in the classes of planar, bipartite, and triangle-free graphs can be equal to the entire set of vertices. So, $\gamma_{[1,2]}$ is not necessarily equal to $\gamma$ in these classes. Etesami et al. (2019) showed that in interval graphs and permutation graphs $\gamma_{[1,3]}(G)=\gamma(G)$. Moreover, it was shown that there exists an interval graph $H$ for which $\gamma(H)<\gamma_{[1,2]}(H)$. Although they proved that for every unit interval graph $G$ we have $\gamma_{[1,2]}(G)=\gamma(G)$, the complexity of finding $\gamma_{[1,2]}$ in non-proper interval graphs remains unsolved.
Another open question in Chellali et al. (2013) was in which graph classes the $[1,j]$-dominating set problem is efficiently solvable. On the negative side, it was also proven that the problem is $NP$-complete even for bipartite graphsChellali et al. (2013). In  Bishnu et al. (2016), the $[1,j]$-dominating sets problem was shown to be $NP$-hard even for chordal and planar graphs. In Etesami et al. (2019) , the authors showed the complexity of the decision problem of whether the $\gamma_{[1,j]}(G)=\gamma(G)$ is an $NP$ problem.
@@
On the positive side, a linear-time algorithms proposed for computing $\gamma_{[1,2]}(G)$ when $G$ is a tree Goharshady et al. (2016), a block graph Sharifani et al. (2020a) and a series-parallel graph Sharifani and Hooshmandasl (2020). For a constant $j$, a polynomial-time algorithm running in roughly $O(n^{j}p(\log n))$ where $p$ is a polynomial function was obtained for $n$-vertex split graphs Bishnu et al. (2016). This is in contrast to the orginal dominating set problem which is NP-hard for this class of graphs. Moreover in Meybodi et al. (2020), a lower bound for computing $[1,j]$-dominating set in the split graph was shown, i.e. there is no algorithm with running time $O(n^{j-\epsilon})$ where $\epsilon>0$ unless $P=NP$. The complexity of $[1,j]$-dominating set problem in various graph classes is summarized in Figure 3. It should be noted that this summary is far from being comprehensive and several other classes of graphs exist. In this paper, we focus only on the circle and interval graphs.
@@
Figure 3: Inclusion relations among well-studied classes of graphs Tripathi et al. (2022) - Complexity of $[1,j]$-dominating set restricted to various classes of graphs in a hierarchy of graphs - $NPc$ indicate that the problem is $NP$-complete, $P$ indicates the problem has a polynomial algorithm, and the question mark indicates the open problem.
@@ 2 Algorithm for interval graphs
Roberts (1969) showed that an interval graph is proper if and only if it contains no induced copy of $K_{1;3}$. It was shown in Dejter (2009) that for any $K_{1;3}$-free graph $\gamma(G)=\gamma_{[1,2]}(G)$. Thus, the following lemma can be easily concluded.
Lemma 1
@@
Another proof for the above lemma was provided in Etesami et al. (2019). The authors also showed that for every interval graph $G$, $\gamma(G)=\gamma_{[1,3]}(G)$ and there exists a graph $G$ such that $\gamma(G)\leq\gamma_{[1,2]}(G)$. However, the complexity of finding a $[1,2]$-dominating set in interval graphs remains open.
In this section, we propose an $O(n^{4})$-time algorithm for computing $\gamma_{[1,2]}(G)$, where $G$ is an interval graph of order $n$. To present and prove the correctness of our algorithm, we need the following lemmas and notation. Let $G=(V,E)$ be an interval graph. Ramalingam and Rangan (1988) proposed a numbering of the vertices of $G$ and stated the following result.
Lemma 2 (Ramalingam and Rangan (1988), Theorem 2.1)
For each interval graph $G=(V,E)$ of order $n$, there is a numbering $(1,2,\ldots,n)$ of vertices in $V$ such that if $ik\in E$ and $i<j<k$, then $jk\in E$. This numbering of vertices can be computed in $O(|V|+|E|)$ time.
In the rest of this section without loss of generality, we assume that $G=(V,E)$ is an interval graph of order $n$ with a numbering $(1,2,\ldots,n)$ of vertices of $G$ satisfying the condition of Lemma 2. Let $a$ and $b$ be integers such that $1\leq a\leq b\leq n$. We introduce the following notation.
@@
    $\gamma^{0}_{[1,2]}(i)=\min\{|D|:D$ is a $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i\notin D\}$,
@@
    $\gamma^{0}_{[1,2]}(j,i)=\min\{|D|:D$ is a $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $x\notin D$ for all $x\in[j,i]\}$,
@@
    $\gamma^{0}_{[1,2]}(j,i:k^{\prime})=\min\{|D|:D$ is a $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $k^{\prime}\in D$ and $x\notin D$ for all $x\in[j,i]\setminus\{k^{\prime}\}\}$, where $k^{\prime}\in[j,i)$,
@@
    $\gamma^{1}_{[1,2]}(i)=\min\{|D|:D$ is a $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i\in D\}$,
@@
    $\gamma^{1}_{[1,2]}(j,i:i)=\min\{|D|:D$ is a $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i\in D$ and $x\notin D$ for all $x\in[j,i)\}$,
@@
    $\gamma^{1}_{[1,2]}(k,i:i,j)=\min\{|D|:D$ is a $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i,j\in D$ and $x\notin D$ for all $x\in[k,i)\setminus\{j\}\}$,
@@
    $\gamma^{1}_{[1,2]}(l,i:i,j,k)=\min\{|D|:D$ is a $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i,j,k\in D$ and $x\notin D$ for all $x\in[l,i)\setminus\{j,k\}\}$,
@@
    $\gamma^{11}_{[1,2]}(l,i:i,j,k)=\min\{|D|:D$ is a $[1,2]$-dominating set of $G[1,i]$ (if exists) such that $i,j,x\in D$ for all $x\in[l,k]$ and $y\notin D$ for all $y\in(k,i)\setminus\{j\}\}$.
@@
Figure 4: (a) A set of intervals corresponding to the vertex set of the interval graph $G$; (b) a numbering $(1,2,\ldots,10)$ of the vertices of $G$ satisfying the condition of Corollary 2; and (c) values $\mathtt{low}(i)$ and $\mathtt{maxlow}(i)$ for all $i\in[1,10]$
Lemma 3 (Poureidi (2022))
@@
Proof 2.1.
It is clear that the case (i) holds. Now, we prove the case (ii). Let $D$ be a $[1,2]$-dominating set of $G[1,i]$ with minimum cardinality, such that $a\notin D$ for all $a\in[i^{\prime},i]$. Thus, $|D|=\gamma^{0}_{[1,2]}(i^{\prime},i)$. Since $i\notin D$, at least one, and at most two of vertices adjacent to $i$ are in $D$. If $i^{\prime}\leq\mathtt{low}(i)$, then there is no vertex in $D$ to dominate $i$, that is, $D$ is not an $[1,2]$-dominating set of $G[1,i]$. Hence, $\mathtt{low}(i)<i^{\prime}$. By case (ii) of Corollary 3, there is a vertex $x\in[\mathtt{maxlow}(i),i]$ such that $xy\notin E$ for each vertex $y\in[1,\mathtt{maxlow}(i))$. Therefore, if $a\notin D$ for all $a\in[\mathtt{maxlow}(i),i]$, then $D$ is not a $[1,2]$-dominating set of $G[1,i]$ and so at least one of vertices of $[\mathtt{maxlow}(i),i)$ is in $D$. Since $a\notin D$ for all $a\in[i^{\prime},i]$, we obtain that $\mathtt{maxlow}(i)<i^{\prime}$ and at least one of vertices of $[\mathtt{maxlow}(i),i^{\prime})$ is in $D$. Assume that $j\in[\mathtt{maxlow}(i),i^{\prime})$ is a vertex in $D$ such that $a\notin D$ for all $a\in(j,i]$. We obtain that $D$ is a $[1,2]$-dominating set of $G[1,j]$. Let $b\in(j,i]$ be a vertex such that $\mathtt{low}(b)\leq\mathtt{low}(a)$ for all $a\in(j,i]$, that is, $\mathtt{low}(b)=\mathtt{low}(j+1,i)$. See Figure 5. Clearly, $\mathtt{low}(j+1,i)\leq\mathtt{low}(i)$. Because $b\notin D$ and $j\in D$, at most one of vertices of $[\mathtt{low}(b),j)$ is in $D$. Hence, either no vertex of $[\mathtt{low}(b),j)$ is in $D$, see Figure 5(a), or exactly one vertex of $[\mathtt{low}(b),j)$ is in $D$, see Figure 5(b). In the following we consider these two cases.